# A new model for the quantum mechanics of the Hydrogen atom

### 16 Mar 2026 v1

@@ 1.1 Overview
The hydrogen atom system is a fundamental well-studied example of a quantum-mechanical system which played a central role in the development of quantum mechanics, see e.g., [Wybourne, LandauLifshitz, MR2156403, Wulfman]. The common model for the spinless non-relativistic system is given by the Hilbert space $L^{2}(\mathbb{R}^{3})$, together with the densely defined (normalized) Schrödinger operator
@@
To deal with degeneracy of the energy eigenspaces one looks for solutions of the Schrödinger equation that are simultaneously eigenfunctions for the total angular momentum operator $L^{2}$ and for its projection on the $z$-axis $L_{z}$. We will refer to these solutions as physical solutions, and denote the linear span of such solutions for a given $E\in\operatorname{Spec}(H_{phys})$ by $\operatorname{PhysSol}(E)$. These solution spaces lead to several layers of symmetries in which the largest symmetry group is $O(4,2)$. In physics, it is called the dynamical symmetry e.g., see [sym12081323]. The model we just described has several flaws.
@@
The first instance is the usual model of the hydrogen atom system, in which there is a clear symmetry of $SO(3)$ and a larger hidden symmetry [^1] In fact there are different hidden symmetry groups within the dynamical symmetry that are fibers of a nice family, See [MR4460278, MR3827131]. of $O(4,2)$ that is called the dynamical symmetry.
@@ 2.1.1 Differential operators on linear spaces
When $X$ is smooth, $\mathcal{D}(X)$ is generated as an algebra by $\mathcal{O}_{X}$ and $\operatorname{Der}(\mathcal{O}_{X})=\{D\in\mathcal{D}^{1}(X)|D(1)=0\}$. In general, beyond the smooth case, $\mathcal{D}(X)$ may not even be Noetherian, e.g., see the case of the cubic cone [cubiccone].
@@ 2.1.2 Differential operators on cones
The following result which follows from Proposition 1.6 of [SSP], explains how a differential operator on the cone can be described by differential operators on the ambient vector space.
@@ 2.2.2 The canonical generating Lie subalgebra
Theorem (Compare with Prop. 3.8 and Cor. 4.4 of [LevasseurAndStafford2017]).
@@
For an explicit embedding of $\mathfrak{so}(n+2,\mathbb{C})$ in $\mathcal{D}(C)$ see section 9 of [LevasseurAndStafford2017] and Theorem 8.2.1. Further details can be found in [Levasseur1986, LevasseurSmithStafford1988, Goncharov1982] and section 1.1 of [KobayashiMano2011].
@@
By direct calculations we see that $\Psi^{Q}(V^{*})\subseteq\mathcal{D}^{2,-1}(V)_{I}$ and for every $w\in V^{*}$, $\Phi^{Q}\circ\Psi^{Q}(w)=(2-n)w$. In addition it is easy to show (and can be deduced from [Goncharov1982]) that $\operatorname{dim}(\mathcal{D}^{2,-1}(V)_{I})\leq\operatorname{dim}(V)$. ∎
@@ 2.2.4 The symmetry group
We remark that a description of $\mathcal{D}(C)$ is essentially given in [Goncharov1982]. We will also give an explicit description in the following section.
@@ 3.0.2 The reductive dual pair symmetry of a PLQS
Howe introduced the notion of a reductive dual pair, see e.g., [MR789290]. Given a reductive Lie algebra $\mathfrak{g}$ and a pair of reductive Lie subalgebras $(\mathfrak{k},\mathfrak{l})$ of $\mathfrak{g}$, the pair is called a reductive dual pair if $\mathfrak{k}$ is equal to the centralizer of $\mathfrak{l}$ in $\mathfrak{g}$ and $\mathfrak{l}$ is equal to the centralizer of $\mathfrak{k}$ in $\mathfrak{g}$.
@@ 5.1 Action of a *-algebra on a Hilbert space
In this section we recall the notion of a partial action of a *-algebra on a Hilbert space and discuss some of its properties. We essentially follow [MR283580] (also see [MR333743, MR3722559]).
@@ 5.1.3 Symmetric, self-adjoint and essentially self-adjoint actions
Definition (Compare with Sec. II &\& IV of [MR283580]).
@@ 5.1.4 An important example of a self-adjoint action
We will prove this claim in Appendix 8.7. Further information on the underlying smooth space of a representation of a Lie group can be found in [BK2014].
@@ 5.3.1 A canonical extension of the representation of $O(\mathcal{V})$ on $\boldsymbol{H}$
By Mackey theory this unitary representation of $P$ is irreducible, also see [KOBAYASHI2003551, Prop. 3.3].
The extension of the representation to $G$: The extension of $\pi_{n}=\pi:{P}\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow\mathcal{U}\left(\boldsymbol{H})\right)$ to a representation of $G$ is based on the fact that $G$ has a distinguished representation $\widetilde{\pi}$ of minimal Gelfand-Kirillov dimension, namely the local theta lift of the trivial representation of $SL_{2}(\mathbb{R})$, see [MR1457244]. Further info on the theta correspondence can be found in [MR546602, MR985172, MR986027].
@@
The hard part of the proof is to show the existence of such an extension. This is proven in [KOBAYASHI2003551, Thm. 3]. ∎
@@
2. For $n>4$ the representation $\widetilde{\pi}_{n}$ is the unique minimal representation of $G$, see [KobayashiMano2011].
3. In this paper we shall only use the Schrödinger model of $\widetilde{\pi}$ which is realized on the canonical Hilbert space $\boldsymbol{H}$. Other important realizations of $\widetilde{\pi}$ exist. Several of them are constructed in [KobayashiMano2011]. Also see [MR2020550, MR1108044, MR1172839].
@@ 5.3.2 The smooth subspace of $\pi$ as a self-adjoint extension
The first part of the proposition is proven in [KobayashiMano2011, Sec. 2.5]. From the explicit formulas by which $\operatorname{Lie}(G)$ acts on $\boldsymbol{H}^{\infty}$ via differential operators on $C_{0}(\mathbb{R})$ e.g. see [KobayashiMano2011, Sec. 2.3], it is evident that the self-adjoint action $\pi^{\infty}:\mathcal{U}(\mathfrak{g})\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow\operatorname{End}(\boldsymbol{H}^{\infty})$ where $\mathfrak{g}$ is the complexification of the Lie algebra $\operatorname{Lie}(G)$ (as in Subsection 5.1.4 ) factors through a quotient surjective map $j:\mathcal{U}(\mathfrak{g})\twoheadrightarrow\mathcal{D}(C)$. As a result we obtain an action of $\mathcal{D}(C)$ on $\boldsymbol{H}^{\infty}$ and in particular $\boldsymbol{H}^{\infty}$ is invariant under the action $\rho^{\infty}$. We denote the obtained action on $\boldsymbol{H}^{\infty}$ by $\rho$.
@@
2. The decomposition $\boldsymbol{H}=\boldsymbol{H}_{+}\oplus\boldsymbol{H}_{-}$ is stable under the connected component of the identity $G_{0}\subset G$, see [KobayashiMano2011, p.19]. This implies that $\boldsymbol{H}^{\infty}=\boldsymbol{H}^{\infty}_{+}\oplus\boldsymbol{H}^{\infty}_{-}$, where $\boldsymbol{H}^{\infty}_{\pm}:=\boldsymbol{H}_{\pm}\cap\boldsymbol{H}^{\infty}$ and moreover $\boldsymbol{H}^{\infty}_{\pm}$ are stable under $\rho$.
@@ 7.1.1 The spectrum of a family of operators
In functional analysis, given a continuous operator $T$ on a dense linear subspace $\boldsymbol{W}$ of a topological vector space $\boldsymbol{H}$ (typically $\boldsymbol{H}$ is a Hilbert space) the spectrum of $T$ is defined as the collection of all $\lambda\in\mathbb{C}$ for which the operator $T-\lambda:\boldsymbol{W}\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow\boldsymbol{H}$ does not have a continuous inverse, see e.g., Section VI3 of [reed]. In physics, given an operator $T$ as above, one is usually interested in a spectral decomposition of $\boldsymbol{H}$ with respect to $T$. Sometimes there are not enough eigenvectors in $\boldsymbol{H}$ for spectral decomposition and a natural extension is the dual space $\boldsymbol{W}^{*}$. This motivates the following definition.
@@ 7.2.3 The $(SL_{2}(\mathbb{R})\times SO(3))$-isotypic decomposition of $\boldsymbol{H}_{\pm}$
This follows from Sections 1.3 and 3.1 of [MR2401813].
@@ 7.2.5 Implications of the Casselman-Wallach theorem
By Casselman-Wallach globalization theorem [Wallach3, Casselman, Wallach2], any two admissible smooth moderate growth Fréchet representations of the same reductive Lie group that are infinitesimally equivalent must be isomorphic as topological representations.
@@ 8.3 A Lie-theoretic description of $\Psi^{Q}$
The Jacobson-Morozov theorem (see e.g., [MR1920389, Ch. X]) for the finite-dimensional simple Lie algebra $\mathfrak{s}$ implies that for $w$ as above there is a unique $f_{w}\in\mathfrak{s}$ such that $\{e_{w},h,f_{w}\}$ is an $\mathfrak{sl}_{2}$-triple in $\mathfrak{s}$ satisfying the standard commutation relations. This uniquely determines the morphism $\psi^{Q}$.
@@ 8.5 Concrete realization of $(\boldsymbol{H}^{\infty}_{\pm})_{\ell}$
The proof immediately follows form the description of the smooth subspace of the discrete series representations of $SL_{2}(\mathbb{R})$ in the Kirillov model as given in [Baruch]. ∎
@@ 8.7 The self-adjoint property of the subspace of smooth vectors
From the definition of the adjoint it is clear that $(\boldsymbol{H}^{\infty})^{\dagger}=\cap_{n}\boldsymbol{H}^{n}$. Note that for $n>m$ the inclusion $\boldsymbol{H}^{n}\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow\boldsymbol{H}^{m}$ is a continuous $G$-equivariant map. Hence it is enough to show that $\boldsymbol{H}^{\infty}=\cap_{n}\boldsymbol{H}^{n}$. Clearly the inclusion $\boldsymbol{H}^{\infty}\subseteq\cap_{n}\boldsymbol{H}^{n}$ is a topological embedding. We will prove the required equality using the fact the $\boldsymbol{H}^{\infty}$ is complete with respect to the system of Sobolev norms $\tau_{n}$. This follows from [BK2014, Lemma. 2.15] (also see [MR84713]).

### 14 Apr 2026 v2

@@ 1.1 Overview
The hydrogen atom system is a fundamental well-studied example of a quantum-mechanical system which played a central role in the development of quantum mechanics, see e.g., [undeff, undefs, undefad, undefai]. The common model for the spinless non-relativistic system is given by the Hilbert space $L^{2}(\mathbb{R}^{3})$, together with the densely defined (normalized) Schrödinger operator
@@
To deal with degeneracy of the energy eigenspaces one looks for solutions of the Schrödinger equation that are simultaneously eigenfunctions for the total angular momentum operator $L^{2}$ and for its projection on the $z$-axis $L_{z}$. We will refer to these solutions as physical solutions, and denote the linear span of such solutions for a given $E\in\operatorname{Spec}(H_{phys})$ by $\operatorname{PhysSol}(E)$. These solution spaces lead to several layers of symmetries in which the largest symmetry group is $O(4,2)$. In physics, it is called the dynamical symmetry e.g., see [undefw]. The model we just described has several flaws.
@@
The first instance is the usual model of the hydrogen atom system, in which there is a clear symmetry of $SO(3)$ and a larger hidden symmetry [^1] In fact there are different hidden symmetry groups within the dynamical symmetry that are fibers of a nice family, See [undefh, undefaf]. of $O(4,2)$ that is called the dynamical symmetry.
@@
Recently we discovered works by Meng [undefx, undefy] that introduce a model close to our model. Meng considers a more general case that includes the case of the Lorentz group acting on a cone.
Meng starts from a simple Euclidean Jordan algebra $J$ and builds a Lie algebra $\mathfrak{co}$, the conformal Lie algebra of $J$, using the Kantor–Koecher–Tits construction.
Then Meng constructs a Kepler cone $\mathcal{C}$ associated to $J$ and considers the Hilbert space $\mathcal{H}=L^{2}(\mathcal{C})$. He observes that the Lie algebra $\mathfrak{co}$ acts on functions on $\mathcal{C}$ via differential operators.
Then he defines a Schrödinger operator in terms of the Lie algebra $\mathfrak{co}$ and computes its discrete spectrum.
In the case of a Jordan algebra $J$ corresponding to Lorentzian space the Kepler cone $\mathcal{C}$ constructed by Meng coincides with our cone $C_{+}(\mathbb{R})$, the Hilbert space $\mathcal{H}=L^{2}(\mathcal{C})$ is our space $\boldsymbol{H}_{+}=L^{2}(C_{+}(\mathbb{R}))$ and his Lie algebra $\mathfrak{co}$ is isomorphic to (a real form of) our quantum dynamical symmetry $\mathfrak{s}$ (see subsection 2.2.2).
Our approach is much more algebraic. We mostly work with the algebra $\mathcal{D}(C)$ of algebraic differential operators on $C$; the Lie algebra $\mathfrak{s}$ arises as a convenient generating subspace of $\mathcal{D}(C)$ and our Schrödinger family lives inside $\mathcal{D}(C)$.
Also, one of the main ingredients of our model is the Schwartz space $\boldsymbol{H}^{\infty}\subset\boldsymbol{H}$ and a self-adjoint action of the algebra $\mathcal{D}(C)$ on the Schwartz space.
The Schwartz space implicitly includes boundary conditions. This allows us to investigate the continuous spectrum in addition to the discrete spectrum.
@@ 2.1.1 Differential operators on linear spaces
When $X$ is smooth, $\mathcal{D}(X)$ is generated as an algebra by $\mathcal{O}_{X}$ and $\operatorname{Der}(\mathcal{O}_{X})=\{D\in\mathcal{D}^{1}(X)|D(1)=0\}$. In general, beyond the smooth case, $\mathcal{D}(X)$ may not even be Noetherian, e.g., see the case of the cubic cone [undefa].
@@ 2.1.2 Differential operators on cones
The following result which follows from Proposition 1.6 of [undefae], explains how a differential operator on the cone can be described by differential operators on the ambient vector space.
@@ 2.2.2 The canonical generating Lie subalgebra
Theorem (Compare with Prop. 3.8 and Cor. 4.4 of [undefu]).
@@
For an explicit embedding of $\mathfrak{so}(n+2,\mathbb{C})$ in $\mathcal{D}(C)$ see section 9 of [undefu] and Theorem 8.2.1. Further details can be found in [undefv, undeft, undefg] and section 1.1 of [undefp].
@@
By direct calculations we see that $\Psi^{Q}(V^{*})\subseteq\mathcal{D}^{2,-1}(V)_{I}$ and for every $w\in V^{*}$, $\Phi^{Q}\circ\Psi^{Q}(w)=(2-n)w$. In addition it is easy to show (and can be deduced from [undefg]) that $\operatorname{dim}(\mathcal{D}^{2,-1}(V)_{I})\leq\operatorname{dim}(V)$. ∎
@@ 2.2.4 The symmetry group
We remark that a description of $\mathcal{D}(C)$ is essentially given in [undefg]. We will also give an explicit description in the following section.
@@ 3.0.2 The reductive dual pair symmetry of a PLQS
Howe introduced the notion of a reductive dual pair, see e.g., [undefj]. Given a reductive Lie algebra $\mathfrak{g}$ and a pair of reductive Lie subalgebras $(\mathfrak{k},\mathfrak{l})$ of $\mathfrak{g}$, the pair is called a reductive dual pair if $\mathfrak{k}$ is equal to the centralizer of $\mathfrak{l}$ in $\mathfrak{g}$ and $\mathfrak{l}$ is equal to the centralizer of $\mathfrak{k}$ in $\mathfrak{g}$.
@@ 5.1 Action of a *-algebra on a Hilbert space
In this section we recall the notion of a partial action of a *-algebra on a Hilbert space and discuss some of its properties. We essentially follow [undefaa] (also see [undefab, undefz]).
@@ 5.1.3 Symmetric, self-adjoint and essentially self-adjoint actions
Definition (Compare with Sec. II &\& IV of [undefaa]).
@@ 5.1.4 An important example of a self-adjoint action
We will prove this claim in Appendix 8.7. Further information on the underlying smooth space of a representation of a Lie group can be found in [undefb].
@@ 5.3.1 A canonical extension of the representation of $O(\mathcal{V})$ on $\boldsymbol{H}$
By Mackey theory this unitary representation of $P$ is irreducible, also see [undefr, Prop. 3.3].
The extension of the representation to $G$: The extension of $\pi_{n}=\pi:{P}\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow\mathcal{U}\left(\boldsymbol{H})\right)$ to a representation of $G$ is based on the fact that $G$ has a distinguished representation $\widetilde{\pi}$ of minimal Gelfand-Kirillov dimension, namely the local theta lift of the trivial representation of $SL_{2}(\mathbb{R})$, see [undefaj]. Further info on the theta correspondence can be found in [undefi, undefl, undefk].
@@
The hard part of the proof is to show the existence of such an extension. This is proven in [undefr, Thm. 3]. ∎
@@
2. For $n>4$ the representation $\widetilde{\pi}_{n}$ is the unique minimal representation of $G$, see [undefp].
3. In this paper we shall only use the Schrödinger model of $\widetilde{\pi}$ which is realized on the canonical Hilbert space $\boldsymbol{H}$. Other important realizations of $\widetilde{\pi}$ exist. Several of them are constructed in [undefp]. Also see [undefq, undefc, undefm].
@@ 5.3.2 The smooth subspace of $\pi$ as a self-adjoint extension
The first part of the proposition is proven in [undefp, Sec. 2.5]. From the explicit formulas by which $\operatorname{Lie}(G)$ acts on $\boldsymbol{H}^{\infty}$ via differential operators on $C_{0}(\mathbb{R})$ e.g. see [undefp, Sec. 2.3], it is evident that the self-adjoint action $\pi^{\infty}:\mathcal{U}(\mathfrak{g})\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow\operatorname{End}(\boldsymbol{H}^{\infty})$ where $\mathfrak{g}$ is the complexification of the Lie algebra $\operatorname{Lie}(G)$ (as in Subsection 5.1.4 ) factors through a quotient surjective map $j:\mathcal{U}(\mathfrak{g})\twoheadrightarrow\mathcal{D}(C)$. As a result we obtain an action of $\mathcal{D}(C)$ on $\boldsymbol{H}^{\infty}$ and in particular $\boldsymbol{H}^{\infty}$ is invariant under the action $\rho^{\infty}$. We denote the obtained action on $\boldsymbol{H}^{\infty}$ by $\rho$.
@@
2. The decomposition $\boldsymbol{H}=\boldsymbol{H}_{+}\oplus\boldsymbol{H}_{-}$ is stable under the connected component of the identity $G_{0}\subset G$, see [undefp, p.19]. This implies that $\boldsymbol{H}^{\infty}=\boldsymbol{H}^{\infty}_{+}\oplus\boldsymbol{H}^{\infty}_{-}$, where $\boldsymbol{H}^{\infty}_{\pm}:=\boldsymbol{H}_{\pm}\cap\boldsymbol{H}^{\infty}$ and moreover $\boldsymbol{H}^{\infty}_{\pm}$ are stable under $\rho$.
@@ 7.1.1 The spectrum of a family of operators
In functional analysis, given a continuous operator $T$ on a dense linear subspace $\boldsymbol{W}$ of a topological vector space $\boldsymbol{H}$ (typically $\boldsymbol{H}$ is a Hilbert space) the spectrum of $T$ is defined as the collection of all $\lambda\in\mathbb{C}$ for which the operator $T-\lambda:\boldsymbol{W}\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow\boldsymbol{H}$ does not have a continuous inverse, see e.g., Section VI3 of [undefac]. In physics, given an operator $T$ as above, one is usually interested in a spectral decomposition of $\boldsymbol{H}$ with respect to $T$. Sometimes there are not enough eigenvectors in $\boldsymbol{H}$ for spectral decomposition and a natural extension is the dual space $\boldsymbol{W}^{*}$. This motivates the following definition.
@@ 7.2.3 The $(SL_{2}(\mathbb{R})\times SO(3))$-isotypic decomposition of $\boldsymbol{H}_{\pm}$
This follows from Sections 1.3 and 3.1 of [undefo].
@@ 7.2.5 Implications of the Casselman-Wallach theorem
By Casselman-Wallach globalization theorem [undefag, undefe, undefah], any two admissible smooth moderate growth Fréchet representations of the same reductive Lie group that are infinitesimally equivalent must be isomorphic as topological representations.
@@ 8.3 A Lie-theoretic description of $\Psi^{Q}$
The Jacobson-Morozov theorem (see e.g., [undefn, Ch. X]) for the finite-dimensional simple Lie algebra $\mathfrak{s}$ implies that for $w$ as above there is a unique $f_{w}\in\mathfrak{s}$ such that $\{e_{w},h,f_{w}\}$ is an $\mathfrak{sl}_{2}$-triple in $\mathfrak{s}$ satisfying the standard commutation relations. This uniquely determines the morphism $\psi^{Q}$.
@@ 8.5 Concrete realization of $(\boldsymbol{H}^{\infty}_{\pm})_{\ell}$
The proof immediately follows form the description of the smooth subspace of the discrete series representations of $SL_{2}(\mathbb{R})$ in the Kirillov model as given in [undef]. ∎
@@ 8.7 The self-adjoint property of the subspace of smooth vectors
From the definition of the adjoint it is clear that $(\boldsymbol{H}^{\infty})^{\dagger}=\cap_{n}\boldsymbol{H}^{n}$. Note that for $n>m$ the inclusion $\boldsymbol{H}^{n}\mathrel{\smash{\meno}}\mathrel{\mkern-3.0mu}\rightarrow\boldsymbol{H}^{m}$ is a continuous $G$-equivariant map. Hence it is enough to show that $\boldsymbol{H}^{\infty}=\cap_{n}\boldsymbol{H}^{n}$. Clearly the inclusion $\boldsymbol{H}^{\infty}\subseteq\cap_{n}\boldsymbol{H}^{n}$ is a topological embedding. We will prove the required equality using the fact the $\boldsymbol{H}^{\infty}$ is complete with respect to the system of Sobolev norms $\tau_{n}$. This follows from [undefb, Lemma. 2.15] (also see [undefd]).